OpenStudy (anonymous):

the difference between 2 #'s is 9. 3 times the larger # is 6 times the smaller. WHAT ARE THE #'S?

OpenStudy (anonymous):

18 and -9

OpenStudy (anonymous):

-9 and 18

OpenStudy (anonymous):

thanks avadh again

OpenStudy (anonymous):

ur welcome chloebella again

OpenStudy (anonymous):

its just 18 and 9

OpenStudy (anonymous):

no negative sign

OpenStudy (anonymous):

thanks cubswinagain

OpenStudy (anonymous):

no.there has to be a negative sign

OpenStudy (anonymous):

nope the difference between 18 and -9 is 27

OpenStudy (anonymous):

Given: x - y = 9 3x = 6y Solve for y: -y = 9 - x Subtract x from both sides y = -9 + x multiply by -1 y=x - 9 clean it up Substitute y: 3x = 6 * (x - 9) multiply 6 into x and -9 Solve for X: 3x = 6x - 54 subtract 6x from both sides 3x - 6x = -54 -3x = -54 Divide by -3 x = 18 Substitute X: y = 9 - 18 y = 9 We get: x = 18 and y = 9 18 - 9 = 9 3(18) = 6(9)